OpenStudy (anonymous):

Need help with calculus problem involving the second fundamental theorem of calculus

OpenStudy (anonymous):

\[\int\limits_{1}^{x^2}(1/t) dt\]

OpenStudy (anonymous):

F(x) = 1 and im supposed to evaluate the problem above

OpenStudy (anonymous):

my problem is that I know for the second fundamental theorem of calc im supposed to replace t with the x^2 but then i still need to integrate so do i integrate from 1 to u?

OpenStudy (anonymous):

\[\int\limits_{1}^{u} (1/x^2)2x dx\]

OpenStudy (anonymous):

is that correct or do i still keep the x^2 on the top?

OpenStudy (superdavesuper):

can u provide a scan of the prob? seems like something is wrong or missing from what u posted...

OpenStudy (anonymous):

sorry i dont have a scan but the problem is exactly like that

OpenStudy (anonymous):

that integral at the top is the same as F(x) which also equals 1

OpenStudy (superdavesuper):

is it F(x)=1 or F(1)=1?

OpenStudy (anonymous):

F(x) = 1

OpenStudy (superdavesuper):

oh.....so the question is asking u to solve for x then?

OpenStudy (anonymous):

im assuming so because it just says to solve the integral

OpenStudy (superdavesuper):

hahahaa its kinda confusing....i thought it was to find the integral or something like that. lets assume it is asking to solve for x okay?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

would it be a good idea to integrate first and then plug in x^2 instead of switching x^2 and t?

OpenStudy (superdavesuper):

Correct @YadielG

OpenStudy (superdavesuper):

do the integral then plug in x^2 and 1 to evaluate the integral.

OpenStudy (superdavesuper):

let \[G(t) = \int\limits_{}^{}\frac{ 1 }{ t}dt,\] F(x) = 1 = G(x^2) - G(1) Then solve for x.

OpenStudy (anonymous):

is there a good rule of thumb for when to actually use the second fundamental theorem of calc? like only use it for definite integrals where the solution is not given

OpenStudy (superdavesuper):

@YadielG in this case, ASSUMING the prob is to solve for x, u dont need the second fundamental theorem as u are not trying to find F'(x).

OpenStudy (anonymous):

That makes sense, maybe im just over thinking things lol

OpenStudy (anonymous):

thank you for your help

OpenStudy (superdavesuper):

err...the operation is to integrate...not differentiate on the 1/t part....so G(t) is not -1/t^2

OpenStudy (irishboy123):

super dave.... thank you☘ \(\large \int\limits_{1}^{x^2} dt \quad (\frac{1}{t} )\) \(\large= \left[\ln t\right]_{1}^{x^2} = 2 \ln x\)

OpenStudy (superdavesuper):

@IrishBoy123 welcome @YadelG let us know if u need more help to solve for x. good luck :)